I NEED ABSOLUTELY CORRECT ANSWERS FOR THESE.Question 1 0.2 pts You receive a credit card offer with terms of 1.25% interest. When you read the fine print, however, you s
Answer #1

1: APR = Periodic rate * Number of periods in a year

= 1.25%*4 = 5%

Answer =5

2: Quarterly interest = 1442-1033 = 409

Quarterly rate = 409/1033 = 39.59%

Effective annual rate =(1+APR/m)^m-1

= (1+39.59%)^4-1

=279.68%

Answer = 279.68